OpenStudy (anonymous):

A fisherman travels downstream at full speed to his favorite fishing spot. The stream is running at a rate of 5 miles per hour, and the trip takes 2 hours to get to the fishing hole. He is surprised, however, to find that the trip takes 6 hours at full speed to get back to the dock when he is finished. How fast will his fishing boat go?

OpenStudy (ankit042):

Let us assume the speed of boat be x miles per hour. For the case of downstream actual speed will be x+5 miles per hour Distance travelled for downstream = 2*(x+5) Effective speed when going upstream will be x-5 Distance for upstream = 6*(x-5) Because two distances are equal we get the equation 2*(x+5) = 6*(x-5) 2x+10 = 6x-30 40 = 4x x = 10 miles per hour